A machine lifts a 40kg mass to a height of 6,000cm in 3s.calculate the power developed by the machine
(Take g=10m/s2
1
Expert's answer
2019-03-11T12:06:54-0400

We can find the power developed by the machine from the formula:


"P = \dfrac{W}{t} = \dfrac{mgh}{t} = \dfrac{40 kg \cdot 10 \dfrac{m}{s^2} \cdot 6000 cm \cdot \dfrac{1 m}{100 cm}}{3 s} = 8000 W."

Answer:

"P = 8000 W."